The length, L, of a rectangle is increased by 50% and the width, W, is doubled to
form a larger rectangle with an area of 30 cm2. What is the largest possible perimeter of the larger rectangle if L and W are integers with L > W?

Answer by richwmiller(17219)

1.50L*2w=30
l>w
original length and width are (5 and 2) or (10 and 1)
testing 5 and 2
7.5 and 4
7.5*2+2*4=15+8=23 cm perimeter
and 10 and 1
15 and 2
15*2+2*2=34 cm perimeter
34> 23 so 34 cm is the largest possible perimeter with integer sides in the original.